OpenStudy (anonymous):

have a 700ml solution of 30% acid, needs a 50% solution of acid. How many ml of 100% acid should be added?

OpenStudy (anonymous):

hmmm ok. so we're saying that the entire solution needs to be .5 (50%)

OpenStudy (anonymous):

so you can solve an equation like this: (.3*700 + 1*x)/(x+700) = .5

OpenStudy (anonymous):

so multiplying both sides by x+700, we get : .5x + 350 = x + 210

OpenStudy (anonymous):

subtract .5 from both sides, and you get .5x + 210 = 350

OpenStudy (anonymous):

so .5x = 140

OpenStudy (anonymous):

140/.5 = x = 280

OpenStudy (anonymous):

so I'd say 280ml

OpenStudy (anonymous):

basically - .5 is the entire fraction of acid in the total solution, hence my equation in the first step - so we know the first part of the end solution since it's given

OpenStudy (anonymous):

700ml * .3 is pure acid

OpenStudy (anonymous):

what you add to the solution, is also pure 100% acid, so 1*x

OpenStudy (anonymous):

the denominator in this case will be the total amount of solution (so the original 700ml plus whatever you added)

OpenStudy (anonymous):

and we know that the entire thing - the fraction - is .5

OpenStudy (anonymous):

and following that logic the answer is 280ml

OpenStudy (anonymous):

i.e. .5 = total ml acid / total ml solution (after you've added everything)

OpenStudy (anonymous):

good luck!

OpenStudy (anonymous):

Thank you